# Flat Spectrum Radio Quasars as high-energy neutrino sources

### v1

@@ 3.2 Radio-quiet / neutrino
Among the 77339 RQ QSOs, 178 fall within the 15 neutrino error regions. Since these numbers do not allow the assesment of any possible significant correlation, we followed two different approaches. First, we compare the radio-quiet QSOs to the FSRQ sample with similar statistic, using the 77339 RQ sources to construct 10,000 different random samples of 1,122 radio-quiet objects each. For each sample, we repeat the same analysis performed on the CLASH FSRQs, extracting subsamples of the same size as the CLASH sample at different flux density limits, ranging from 1122 to 270 objects. For each of these samples we count the matches with the 15 neutrino positions and estimated the expected number of chance coincidences again by randomly shifting the positions. We then repeated the same analysis on the 1,122 most massive and the 1,122 most luminous QSOs, regardless of their radio flux.
@@
The p-value results are shown in Fig. 4. We find average p-values of 0.7 with the 99% of the samples having p-values higher than 0.02, meaning that the number of RQ QSOs present in the neutrino error box is fully consistent with what expected from chance coincidence.
@@
The absence of a correlation between radio-quiet QSOs and neutrino events, combined with the correlation found for the FSRQ subsample, strongly suggests that, if FSRQs are confirmed as neutrino sources, the production mechanism must be tied to the physics of the jet rather than the accretion flow.
@@ 4 Neutrinos and the closest CLASH FSRQs
In the previous section, we assessed the correlation between neutrinos and CLASH objects by counting the number of neutrino events that include FSRQs within their 90% confidence error regions. In this Section, we use a different approach by analyzing the angular distances between each of the 15 considered neutrino events and the nearest CLASH source. Fig. 5 presents the minimum distances between the 15 neutrino positions and CLASH sources as a function of declination, considering only objects with a flux density above 100 mJy (497 sources). The figure also includes the mean and standard deviation of the distribution of minimum distances from a random sample for comparison. A clear trend emerges when comparing the subsamples at low and high declinations. At low declinations, the distances between 7 out of 8 neutrino events and the nearest CLASH source are consistent with random expectations. In contrast, at higher declinations, all 7 neutrino events have a nearby CLASH bright object within 0.7∘, significantly closer than expected from random distributions.
@@
We apply a Kolmogorov-Smirnov (KS) test to evaluate whether the CLASH-neutrino distance distribution differs significantly from a random distribution. The results, for both the full sample of 15 neutrino events and the high-declination subsample (7 events), are presented in Fig. 4 and summarized in Tab. 2 for different radio flux density thresholds. While the difference is moderately significant for the full sample, the high-declination subsample shows a much stronger deviation from randomness. The minimum uncorrected p-value at flux limit of 100 mJy is 410-6, corresponding to a significance of 4.0 after accounting for the six trials (Fig. 4).
@@
The dependence of the distance between the closest CLASH-FSRQ and neutrino events on declination may have a technical explanation. The performance of the IceCube telescope, particularly in terms of background levels and effective area, varies significantly with declination (Abbasi et al. 2023b), which corresponds to the (opposite of the) neutrino incident direction. Although we are not in a position to fully assess the event detection error budget, our results suggest that the positional accuracy and purity level of IceCat-1 catalog events are higher at declinations above 20∘.
@@
As the number of events with good positional accuracy in IceCat-1 within our region of interest is limited, in order to check this result we expand our sample. We consider all the CLASS catalog, regardless of identification with an S11 QSO. In order to be homogeneous with the S11 catalog, we consider only source with magnitude 20 in at least one of the SDSS band. With this limit at flux density 100 mJy, the CLASS catalog includes 1790 objects, among which we find that 1052 are classified as FSRQ in the literature, while the remaining objects are unclassified.
@@
This extension allows us to use all the 33 events in the northern hemisphere with good positional accuracy (as defined earlier). Additionally, we include the 18 neutrino events that meet the same positional accuracy criteria and occurred since January 2022, as reported in the GCN catalog 222 https://gcn.gsfc.nasa.gov/amon_icecube_gold_bronze_events.html. The distances between these 51 events and the closest CLASS source are reported in Fig. 7. Among the four objects added to the high declination sample, three have a minimum distance 0.4∘, while one is at 1.8∘. We note that one of the three at short distance is GB6J0740+28, a CLASH object in the error region of IC200117A which is excluded from the sample of 15 events because too close to the border of our region of interest (Sect. 2.2). Therefore, the pattern of high declination events being closer to a bright radio source seems to be present also in this expanded catalog.
5 High declination sample
@@
Focusing on the original sample of 15 events, we find that, among the seven neutrinos detected at declinations above 20∘, five have a bright CLASH FSRQ (flux 100 mJy) within their 90% confidence error region, while all seven have at least one CLASH FSRQ within 0.7∘ (Fig. 5 ). We note that the two high declination events without a bright CLASH source within their nominal 90% confidence error region would include such a source if a systematic error of 0.1∘ were (linearly) added to their nominal positional uncertainties (Fig. 6 ).
@@
In this section, we highlight the main characteristics of these 7 CLASH sources. All the information presented here is derived from public databases and catalogues. The seven FSRQs have redshift in the 0.5-2.3 range and optical AB magnitude between 17 and 19. X-ray fluxes in the 0.5-2.0 keV band vary between 1 and 5 10-13 erg s-1 cm-2 which are typical values for FSRQ with radio flux 100 mJy. The SMBH masses range between 108.8 and 109.7 solar masses and are accreting at rates estimated between 0.1 and 0.8 in terms of Eddington ratios (Tab. 3). We investigate whether these numbers are consistent with the characteristics of the S11 RQ QSO population. The analysis is hampered by the small sample size and no significant difference was detected, with K-S test probabilities exceeding 10% in all three cases.
From a visual inspection of the SEDs of these 7 objects, based on the FIRMAMENTO333https://firmamento.hosting.nyu.edu/home tool (Giommi 2025), we find that they all have a usual blazar shape, with a synchrotron peak falling at 1013-1014 Hz, i.e. in the typical range observed in FSRQ.
@@
Only one object listed in Tab. 3 is detected in gamma-rays by FERMI, being listed as positionally associated with a source in the 4FGL-DR4 catalogue (Abdollahi et al. 2020; Ballet et al. 2023), while a second one (GB6J1125+2610) is indicated only as possible, low confidence, association. Therefore, only a small fraction of the FSRQ that are likely associated to a neutrino event seem to be clearly detected in gamma-rays. Neutrinos produced via photo-pion interactions are expected to be accompanied by an equivalent emission of -ray photons. However, the neutrino production site is likely optically thick to high-energy photons (e.g. Böttcher, 2019) and, therefore, the gamma-rays produced co-spatially with neutrinos are expected to be strongly absorbed and not detectable. In this scenario, the observed -ray emission can be attributed to the leptonic component through the Inverse Compton mechanism. FSRQ with a low frequency peak (1013-1014 Hz) usually have the high-energy hump peaking at energies 100 MeV, i.e. below the sensitivity of FERMI. These blazars could appear weak in FERMI, even if their luminosity at the peak is high. Notably, the detection of one (possibly two) sources is consistent with the detection rate of bright FSRQs (flux 100 mJy) by Fermi, which stands at approximately 40%.
@@
One of these objects (GB6J1125+2610, aka TXS 1123+264) has been previously indicated as a possible neutrino emitter by Plavin et al. (2020). Interestingly, a study of 20 years of archival VLBI data has detected a significant brightening of the core after the neutrino event (Kőmíves et al. 2024).
@@
Moreover, we compared the positions of the seven FSRQ candidates with the 9-year IceCube sky map (IceCube Collaboration et al., 2022), which provides the probability that the observed neutrino flux is of non-astrophysical origin, mapped across the entire sky on a 0.2 0.2∘ grid. Interestingly, we find that 4 out of the 7 FSRQ candidates lie within 1∘ of a significant peak (local p-value ¡ 410-4) in the sky map, whereas only 0.12 such associations would be expected by chance. The most intriguing case is GB6J1350+233 (CRATESJ1350+23), located approximately 0.8∘ from the fifth most significant position on the IceCube scan (IceCube Collaboration et al., 2022, Tab. S2). Although the sky map and the IceCat-1 catalog are not independent datasets, this correlation might suggest that, with high probability, some of these seven FSRQs produce detectable and measureable neutrino fluxes.
6 Bronze, gold events and the fraction of neutrinos produced by FSRQ
@@
In the IceCat-1 catalog, neutrino events are classified as either bronze or gold, based on their estimated probability of being genuine astrophysical in origin. Bronze events are assigned a 30% probability, while gold events have a 50% probability (Abbasi et al., 2023b). In the sample of 15 neutrino events considered in our analysis, six are flagged as gold and nine as bronze, leading to an expected total of 5.7 real astrophysical neutrino events. If we define an association as a positional match within 0.7∘, we have that 8 out of 15 neutrinos have at least one associated CLASH FSRQ, with one event (IC180125A) having two associations. Using our procedure to generate control samples with randomized QSO positions, we estimate that only 1.5 such associations are expected by chance. This implies 7–8 genuine associations, which is consistent with the expected 5.7 astrophysical, at least under the hypothesis that a significant fraction of them are produced by FSRQs. However, focusing only on the high-declination subsample yields a different picture. As noted, all seven neutrino events located at declinations above 20∘ are associated with a CLASH FSRQ, compared to only 0.8 associations expected by chance. Among these, five events are bronze and two are gold, implying an expected 2.5 real astrophysical events, inconsistent with the observed 7 associations (probability 1%). As discussed earlier, the observed difference between the high and low declination samples may be due to unaccounted systematic uncertainties, affecting either the positional accuracy or the purity of the IceCat-1 catalog. The fact that seven out of seven high-declination events are associated with a nearby FSRQ suggests that this subsample may indeed have a higher astrophysical purity.
@@
Assuming that all the 7 neutrino are real, this would imply that the fraction of neutrino produced by FSRQs is 60% at 2 confidence. Such a high fraction of high-energy neutrino events originating from FSRQs appears to be in tension with the much lower values reported in other studies (20% Palladino et al., 2020; Oikonomou, 2022, and references therein). The first argument in those studies is based on the expected electromagnetic counterpart, primarily in the gamma-ray band. These works typically estimate upper limits on the total neutrino flux from blazars by assuming a fixed neutrino/-ray flux ratio. However, as discussed in the previous section, many FSRQs may appear faint or even undetectable by Fermi, as their high-energy spectral peak may fall below the Fermi LAT sensitivity range. As a result, the contribution of such sources to the neutrino flux could be significantly underestimated in gamma-ray-based analyses.
@@
A second argument comes from the lack of significant neutrino point sources, which, combined with the observation of a diffuse astrophysical neutrino flux, implies that the dominant sources should form a dense population of weak emitters, rather than a few luminous ones. In a recent analysis, Abbasi et al. (2023c) found that, assuming steady emission and evolution following the star formation rate (SFR), a minimum local source density of 710-9 Mpc-3 is required to match the observed neutrino flux. FSRQs, with a local density of only 5.610-11 Mpc-3, as inferred from the Fermi catalog, are too rare to contribute significantly. However, this conclusion is highly dependent on the assumed source evolution. In the same study (Abbasi et al., 2023c), the authors note that adopting the evolution of the AGN luminosity function would reduce the required source density by nearly two orders of magnitude, down to 610-11Mpc-3 , much closer to the FSRQ density.
@@
In our analysis, we find that the neutrino-associated FSRQs have radio luminosities at 1.4 GHz in the range of 1042-1044 erg s-1. The space density and redshift evolution of this population have been well characterized in previous works (Mao et al., 2017; Caccianiga et al., 2019; Diana et al., 2022). Their local space density is estimated to be around 410-10 Mpc-3, with a strong positive evolution peaking at z 1.9, where the density reaches 2.5 10-8 Mpc-3. This evolution is significantly faster than that of the SFR and more closely resembles that of AGN. Taking this into account, our findings are consistent with both the absence of neutrino multiplets and the observed diffuse neutrino flux and thus reconcile the high apparent association rate of FSRQs with the broader neutrino source population. A similar result has been found by Neronov & Semikoz (2020).
@@
QSO Name | Other Names | RA | DEC | z | S5GHz | Log Lbol | Log | Edd. ratio | FERMI | IceCat-1 | dist |

### v2

@@ 3.2 RQ quasar–neutrino association
Among the 77339 RQ quasars, 178 fall within the 15 neutrino error regions. Since these numbers do not allow the assessment of any possible significant correlation, we followed two different approaches. First, we compared the RQ quasars to the FSRQ sample with similar statistic, using the 77339 RQ sources to construct 10,000 different random samples of 1,122 RQ objects each. For each sample, we repeated the same analysis performed on the CLASH FSRQs, extracting subsamples of the same size as the CLASH sample at different flux density limits, ranging from 1122 to 270 objects. For each of these samples, we counted the matches with the 15 neutrino positions and estimated the expected number of chance coincidences again by randomly shifting the positions. We then repeated the same analysis on the 1,122 most massive and the 1,122 most luminous quasars, regardless of their radio flux.
@@
The p-value results are shown in Fig. 4. We find average p-values of 0.7 with the 99% of the samples having p-values higher than 0.02, meaning that the number of RQ quasars present in the neutrino error box is fully consistent with what expected from chance coincidence.
@@
The absence of a correlation between RQ quasars and neutrino events, combined with the correlation found for the FSRQ subsample, strongly suggests that, if FSRQs are confirmed as neutrino sources, the production mechanism must be tied to the physics of the jet rather than the accretion flow.
@@ 4 Neutrinos and the closest CLASH FSRQs
In the previous section we assessed the correlation between neutrinos and CLASH objects by counting the number of neutrino events that include FSRQs within their 90% confidence error regions. In this section we present a different approach, analysing the angular distances between each of the 15 considered neutrino events and the nearest CLASH source. Figure 5 presents the minimum distances between the 15 neutrino positions and CLASH sources as a function of declination, considering only objects with a flux density above 100 mJy (497 sources). The figure also shows, for comparison, the mean and standard deviation of the distribution of minimum distances obtained from random control samples. Specifically, we generated 10,000 control samples of 497 elements, randomly extracted from the RQ population, which uniformly covers the same area as the CLASH sample. As an additional check, we produced 10,000 further samples by randomizing the positions within a circular annulus of radius 3–10∘ around the original locations of the CLASH and RQ quasars, finding perfectly consistent numbers.
A clear trend emerges when comparing the subsamples at low and high declinations. At low declinations, the distances between seven out of the eight neutrino events and the nearest CLASH source are consistent with random expectations. In contrast, at higher declinations, all seven neutrino events have a nearby CLASH bright object within 0.7∘, significantly closer than expected from random distributions.
@@
We applied a Kolmogorov-Smirnov (KS) test to evaluate whether the CLASH-neutrino distance distribution differs significantly from a random distribution. The results, for both the full sample of 15 neutrino events and the high-declination subsample (7 events), are presented in Fig. 4 and summarized in Table 2 for different radio flux density thresholds. While the difference is moderately significant for the full sample, the high-declination subsample shows a much stronger deviation from randomness. The minimum uncorrected p-value at flux limit of 100 mJy is 410-6, corresponding to a significance of 4.0 after accounting for the six trials (Fig. 4).
@@
The dependence of the distance between the closest CLASH-FSRQ and neutrino events on declination may have a technical explanation. The performance of the IceCube telescope, particularly in terms of background levels and effective area, varies significantly with declination (Abbasi et al. 2023b), which corresponds to the (opposite of the) neutrino incident direction. Although we are not in a position to fully assess the event detection error budget, our results suggest that the positional accuracy and purity level of IceCat-1 catalogue events are higher at declinations above 20∘.
@@
As the number of events with good positional accuracy in IceCat-1 within our region of interest is limited, in order to check this result, we expanded our sample. We considered the entire CLASS catalogue, regardless of identification with an S11 quasar. To be homogeneous with the S11 catalogue, we considered only sources with magnitudes of 20 in at least one of the SDSS band. With this limit at flux densities of 100 mJy, the CLASS catalogue includes 1790 objects, of which 1052 are classified as FSRQs in the literature; the remaining objects are unclassified.
@@
This extension allowed us to use all the 33 events in the northern hemisphere with good positional accuracy (as defined earlier). Additionally, we included the 18 neutrino events that meet the same positional accuracy criteria and occurred from January 2022 to December 2024, as reported in the GCN catalogue444https://gcn.gsfc.nasa.gov/amon_icecube_gold_bronze_events.html. The distances between these 51 events and the closest CLASS source are reported in Fig. 7. Among the four objects added to the high-declination sample, three have a minimum distance 0.4∘, while one is at 1.8∘. We note that one of the three at a short distance is GB6J0740+28, a CLASH object in the error region of IC200117A, which was excluded from the sample of 15 events because it is too close to the border of our region of interest (Sect. 2.2). Therefore, the pattern of high-declination events being closer to a bright radio source seems to also be present in this expanded catalogue.
5 High-declination sample
@@
Focusing on the original sample of 15 events, we find that, among the seven neutrinos detected at declinations above 20∘, five have a bright CLASH FSRQ (flux 100 mJy) within their 90% confidence error region, while all seven have at least one CLASH FSRQ within 0.7∘ (Fig. 5). We note that the two high-declination events without a bright CLASH source within their nominal 90% confidence error region would include such a source if a systematic error of 0.1∘ were (linearly) added to their nominal positional uncertainties (Fig. 6 ).
@@
In this section we highlight the main characteristics of these seven CLASH sources. All the information presented here is derived from public databases and catalogues. The seven FSRQs have redshifts in the 0.5-2.3 range and optical AB magnitudes between 17 and 19. X-ray fluxes in the 0.5-2.0 keV band vary between 1 and 5 10-13 erg s-1 cm-2, which are typical values for FSRQs with radio fluxes 100 mJy. The SMBH masses range between 108.8 and 109.7 solar masses and are accreting at rates estimated between 0.1 and 0.8 in terms of Eddington ratios (Table 3). We investigated whether these numbers are consistent with the characteristics of the S11 RQ quasar population. The analysis is hampered by the small sample size and no significant difference was detected, with K-S test probabilities exceeding 10% in all three cases.
@@
From a visual inspection of the spectral energy distribution of these seven objects, determined using the FIRMAMENTO555https://firmamento.hosting.nyu.edu/home tool (Giommi 2025), we find that they all have a usual blazar shape, with a synchrotron peak falling at 1013-1014 Hz, i.e. in the typical range observed in FSRQs. Only one object listed in Table 3 was detected in gamma rays by Fermi; it is listed as positionally associated with a source in the 4FGL-DR4 catalogue (Abdollahi et al. 2020; Ballet et al. 2023). A second object (GB6J1125+2610) is indicated only as a possible, low-confidence, association. Therefore, only a small fraction of the FSRQs that are likely associated with a neutrino event seem to be clearly detected in gamma rays. Neutrinos produced via photo-pion interactions are expected to be accompanied by an equivalent emission of -ray photons. However, the neutrino production site is likely optically thick to high-energy photons (e.g. Böttcher, 2019), and therefore, the gamma rays produced co-spatially with neutrinos are expected to be strongly absorbed and not detectable. In this scenario, the observed -ray emission can be attributed to the leptonic component through the inverse Compton mechanism. FSRQs with a low frequency peak (1013-1014 Hz) usually have the high-energy hump peaking at energies 100 MeV, i.e. below the sensitivity of Fermi. These blazars could appear weak in Fermi, even if their luminosity at the peak is high. Notably, the detection of one (possibly two) sources is consistent with the detection rate of bright FSRQs (flux 100 mJy) by Fermi, which stands at approximately 40%.
@@
One of these objects (GB6J1125+2610, also known as TXS 1123+264) has been previously indicated as a possible neutrino emitter by Plavin et al. (2020). Interestingly, a study of 20 years of archival very long-baseline interferometry (VLBI) data detected a significant brightening of the core after the neutrino event (Kőmíves et al. 2024).
@@
Moreover, we compared the positions of the seven FSRQ candidates with the 9-year IceCube sky map (IceCube Collaboration et al., 2022), which provides the probability that the observed neutrino flux is of non-astrophysical origin, mapped across the entire sky on a 0.2 0.2∘ grid. Interestingly, we find that four out of the seven FSRQ candidates lie within 1∘ of a significant peak (local p-value ¡ 410-4) in the sky map To calculate the number of associations expected by chance, we exploited the uniform distribution of RQ quasar positions over the region by extracting 10,000 samples of seven positions each. We find a mean value of 0.12. The most intriguing case is GB6J1350+233 (CRATESJ1350+23), located approximately 0.8∘ from the fifth most significant position on the IceCube scan (IceCube Collaboration et al., 2022, Table S2). Although the sky map and the IceCat-1 catalogue are not independent datasets, this correlation might suggest that, with a high probability, some of these seven FSRQs produce detectable and measurable neutrino fluxes.
6 Bronze and gold events and the fraction of neutrinos produced by FSRQs
@@
In the IceCat-1 catalogue, neutrino events are classified as either bronze or gold, based on their estimated probability of being genuinely astrophysical in origin. Bronze events are assigned a 30% probability, while gold events have a 50% probability (Abbasi et al., 2023b). In the sample of 15 neutrino events considered in our analysis, six are flagged as gold and nine as bronze, leading to an expected total of 5.7 real astrophysical neutrino events. If we define an association as a positional match within 0.7∘, 8 out of the 15 neutrinos have at least one associated CLASH FSRQ, with one event (IC180125A) having two associations. Using our procedure to generate control samples with randomized quasar positions, we estimate that only 1.5 such associations are expected by chance. This implies 7–8 genuine associations, which is consistent with the expected 5.7 astrophysical, at least under the hypothesis that a significant fraction of them are produced by FSRQs. However, focusing only on the high-declination subsample yields a different picture. As noted, all seven neutrino events located at declinations above 20∘ are associated with a CLASH FSRQ, compared to only 0.8 associations expected by chance. Among these, five events are bronze and two are gold, implying an expected 2.5 real astrophysical events, inconsistent with the observed 7 associations (probability 1%). As discussed earlier, the observed difference between the high- and low-declination samples may be due to unaccounted systematic uncertainties, affecting either the positional accuracy or the purity of the IceCat-1 catalogue. The fact that all seven high-declination events are associated with a nearby FSRQ suggests that this subsample has a higher astrophysical purity.
@@
Assuming that all the 7 neutrino are real, this would imply that the fraction of neutrino produced by FSRQs is 60% at 2 confidence. Such a high fraction of high-energy neutrino events originating from FSRQs appears to be in tension with the much lower values reported in other studies (20%, Palladino et al., 2020; Oikonomou, 2022, and references therein). The first argument in those studies is based on the expected electromagnetic counterpart, primarily in the gamma-ray band. These works typically estimated upper limits on the total neutrino flux from blazars by assuming a fixed neutrino/-ray flux ratio. However, as discussed in the previous section, many FSRQs can appear faint or even undetectable by Fermi, as their high-energy spectral peak can fall below the Fermi LAT sensitivity range. As a result, the contribution of such sources to the neutrino flux could be significantly underestimated in gamma-ray-based analyses.
@@
A second argument comes from the lack of significant neutrino point sources, which, combined with the observation of a diffuse astrophysical neutrino flux, implies that the dominant sources should form a dense population of weak emitters, rather than a few luminous ones. In a recent analysis, Abbasi et al. (2023c) found that, assuming steady emission and evolution following the star formation rate, a minimum local source density of 710-9 Mpc-3 is required to match the observed neutrino flux. FSRQs, with a local density of only 5.610-11 Mpc-3, as inferred from the Fermi catalogue, are too rare to contribute significantly. However, this conclusion is highly dependent on the assumed source evolution. In the same study (Abbasi et al., 2023c), the authors note that adopting the evolution of the AGN luminosity function would reduce the required source density by nearly two orders of magnitude, down to 610-11Mpc-3 , much closer to the FSRQ density.
@@
In our analysis we find that the neutrino-associated FSRQs have radio luminosities at 1.4 GHz in the range of 1042-1044 erg s-1. The space density and redshift evolution of this population have been well characterized in previous works (Mao et al., 2017; Caccianiga et al., 2019; Diana et al., 2022). Their local space density is estimated to be around 410-10 Mpc-3, with a strong positive evolution peaking at z 1.9, where the density reaches 2.5 10-8 Mpc-3. This evolution is significantly faster than that of the star formation rate and more closely resembles that of AGNs. Taking this into account, our findings are consistent with both the absence of neutrino multiplets and the observed diffuse neutrino flux and thus reconcile the high apparent association rate of FSRQs with the broader neutrino source population. A similar result has been found by Neronov & Semikoz (2020).
Abbasi et al. (2022b), using a stacking analysis of 137 FSRQs selected in the 30–100 MeV band, found that their contribution to the unresolved background is . At first glance, this result may seem to be in tension with our findings, since some level of soft -ray/hard X-ray emission is inevitably expected from FSRQs if they are indeed neutrino emitters (see the discussion in Abbasi et al. (2022b). However, the expected contribution of these very bright and rare objects to the unresolved neutrino flux is not significant. Furthermore, the hadronic contribution to the soft -ray flux is difficult to quantify, while the inverse Compton emission is certainly contributing in this band. It is therefore plausible that the 137 FSRQs bright in the soft -ray band considered by Abbasi et al. (2022b) are simply those with the strongest Compton dominance, and thus not necessarily the most promising neutrino sources. We do expect that the FSRQs discussed in this work emit in the soft -ray band, but it is likely that the sensitivity of Fermi in this energy range is insufficient to detect them.
@@
| Quasar Name | Other Names | RA | Dec. | z | S5GHz | Log Lbol | Log | Edd. ratio | Fermi | IceCat-1 | dist |